Solve: 7 / 7 – 5 / 7 Solve: 7 / 7 – 5 / 7 Solve: 7 / 7 – 5 / 7

Solve:
  7 / 7 – 5 / 7
Solve:
  7 / 7 – 5 / 7
Solve:
  7 / 7 – 5 / 7

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

7 / 7–5 / 7
= (7–5) / 7
=2 / 7

​Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Here denominator iis same so we need not take the L.C.M.
7/7 - 5/7 = 2/7
 
Hope it helps!
Thank you ✌️